# Xu Shen  and  Yuqiang Zheng Morningside Center of Mathematics Academy of Mathematics and Systems Science Chinese Academy of Sciences

### v2

@@ 2.2. Smooth Pappas-Rapoport splitting models
By abuse of notation, we still denote by $\mathcal{V}$ the integral automorphic vector bundle over $\operatorname{\mathscr{A}}^{\mathrm{spl}}$. We discuss the example of standard representation $\Lambda^{\mathrm{spl}}$. Let $(A^{\textrm{univ}},\lambda,\iota)$ be the universal abelian scheme over $\operatorname{\mathscr{A}}^{\mathrm{spl}}$ and $\mathcal{H}=H^{1}_{\operatorname{dR}}(A^{univ}/\operatorname{\mathscr{A}}^{\mathrm{spl}})$. Then as before we have decomposition $\mathcal{H}=\bigoplus_{i}\mathcal{H}_{i}^{m_{i}},\mathcal{H}_{i}=\bigoplus_{j}\mathcal{H}_{i,j}$. We have also the universal splitting structure $\underline{\operatorname{\mathscr{F}}_{\bullet}}$ on $\mathcal{H}$. For each $i,j,l$, we define
@@ 3.1. $F$-zips and $G$-zips
  • $C^{\bullet}=(C^{i})_{i\in\mathbb{Z}}$ is a descending filtration on $M$ and each $C^{i}$ are locally free $\operatorname{\mathcal{O}}_{S}$-module;
@@
  • $\varphi_{\bullet}=(\varphi_{i})_{i\in\mathbb{Z}}$ and for each $i$, $\varphi_{i}:C^{i}/C^{i+1}\to D_{i}/D_{i-1}$ is a $\sigma$-linear map whose linearizion
@@
Next we briefly review the notion of $G$-zips and the theory of $G$-zip stacks ([51, 52]) for later use in this section. Let $G$ be a connected reductive group over $\mathbb{F}_{p}$, and $\chi$ a cocharacter of $G$ defined over a finite field $\kappa|\mathbb{F}_{p}$. Let $P_{+}$ (resp. $P_{-}$) be the parabolic subgroup of $G_{\kappa}$ such that its Lie algebra is the sum of spaces with non-negative weights (resp. non-negative weights) in $\operatorname{Lie}(G_{\kappa})$ under $\operatorname{Ad}\circ\chi$. We will also write $U_{+}$ (reps. $U_{-}$) for the unipotent radical of $P_{+}$ (resp. $P_{-}$). Let $L$ be the common Levi subgroup of $P_{+}$ and $P_{-}$.
@@ A $G$-zip of type $\chi$ over $S$ is a tuple $\underline{I}=(I,I_{+},I_{-},\iota)$ consisting of
[…]
@@
We have an isomorphism of algebraic stacks $[E_{G,\chi}\backslash G_{\kappa}]\simeq{G}\text{-}\mathrm{Zip}^{\chi}_{\kappa}$.
@@
    Denote by $|[E_{G,\chi}\backslash G_{\kappa}]\otimes k|$ the topological space of $[E_{G,\chi}\backslash G_{\kappa}]\otimes k$, and ${}^{J}W$ the topological space induced by the partial order $\preceq$. Then the association $w\mapsto G_{w}$ induces a homeomorphism ${}^{J}W\simeq|[E_{G,\chi}\backslash G_{\kappa}]|$.
@@
    ¯$\underline{\mathscr{F}_{\bullet}}=(\mathscr{F}_{i,j}^{l})$ is a splitting structure of $\operatorname{\mathcal{O}}_{F}\otimes_{\mathbb{Z}_{p}}\operatorname{\mathcal{O}}_{S}$-module $\omega_{A/S}$. This means: if we write $\mathcal{H}=\mathcal{H}_{\operatorname{dR}}^{1}(A/S)$, it is an $\operatorname{\mathcal{O}}_{F}\otimes\operatorname{\mathcal{O}}_{S}$-module, hence has a decomposition
@@
Note that each $\mathcal{M}_{i,j}^{l}$ is a locally free $\operatorname{\mathcal{O}}_{S}$-module of rank $2d_{i}$. As in the description blew Definition 2.1, in fact we have some reduced description for $\mathcal{M}_{i,j}^{l}$ according to the type of $i$, which we leave to the reader. Now we write
@@
    which is an $\operatorname{\mathcal{O}}_{S}$-linear morphism. On the other side, the injection $\varepsilon_{i}^{-1}\mathscr{F}_{i,j}^{l-2}\hookrightarrow\varepsilon_{i}^{-1}\mathscr{F}_{i,j}^{l-1}$ induce an $\operatorname{\mathcal{O}}_{S}$-linear morphism
@@
[…]
@@
[…]
@@
There is a canonical isomorphism $g_{i}:\operatorname{\mathcal{M}}^{1}_{i}\to\operatorname{\mathcal{M}}^{1}_{i}$ sends $\operatorname{Ker}(V_{i}^{1})$ to $\operatorname{Im}(F_{i}^{1})$.
@@
For each $j$, fix an isomorphism $\operatorname{\mathcal{H}}_{i,j}\simeq R[\varepsilon_{i}]^{2d_{i}}$ and lift it to ~$\widetilde{\operatorname{\mathcal{H}}}_{i,j}=R[[t]]^{2d_{i}}$ such that ~$\widetilde{\operatorname{\mathcal{H}}}_{i,j}/(t^{e_{i}})=\operatorname{\mathcal{H}}_{i,j}$. For each $\operatorname{\mathscr{F}}_{i,j}^{l}\subset\operatorname{\mathcal{H}}_{i,j}$, let ~$\widetilde{\operatorname{\mathscr{F}}}_{i,j}^{l}$ be its lifting in ~$\widetilde{\operatorname{\mathcal{H}}}_{i,j}$. One can lift the Frobenius and Verschiebung morphisms of $\operatorname{\mathcal{H}}_{i}$ to injections of ~$\widetilde{\operatorname{\mathcal{H}}}_{i}$, such that
@@
[…]
Such isomorphism sends the lifting of $\operatorname{Ker}(V_{i,j}^{1})$ to lifting of $\operatorname{Im}(F_{i,j}^{1})$. After mod $(t^{e_{i}})$ of above isomorphism, we get an isomorphism
@@
Such datum gives the parabolic subgroup $P_{+},P_{-}$ of $\operatorname{\mathcal{G}}_{\kappa}^{\mathrm{spl}}$ and corresponding unipotent group $U_{+},U_{-}$, and common Levi subgroup $L=P_{+}\cap P_{-}$ as in previous subsection. The cocharacter $\mu$ induces a standard $F$-zip
@@
and $D^{\phi}$ is the same module $D$ but with $\operatorname{\mathcal{G}}_{0}^{\mathrm{spl}}$-action twist by $\phi$. The sub vector bundles $C$ and $D$ are totally isotropic with respective to the symplectic form $\psi$ on $\operatorname{\mathcal{M}}$.
@@
[…]
where $M^{\mathrm{loc}}$ is the special fiber of $\mathbb{M}^{\mathrm{loc}}(\mathcal{G}^{\mathrm{spl}},\mu)$, $q$ sends $(x=(\underline{A},\underline{\operatorname{\mathscr{F}}}_{\bullet}),\tau:\operatorname{\mathcal{M}}_{x}\simeq\Lambda^{\mathrm{spl}}_{0})$ to $\tau(\operatorname{\mathcal{M}}_{x}\supset C_{x})\in M^{\mathrm{loc}}=\operatorname{\mathcal{G}}_{0}^{\mathrm{spl}}/P_{+}$. By same arguments in [57, §3.4], we have
@@
where $M^{\mathrm{loc},c}(\operatorname{\mathcal{G}}_{0},\mu)$ classifies the conjugate filtration of $\operatorname{\mathscr{A}}_{0}$ defined in [57] and $q^{c}$ is a smooth morphism (same as [57, Theorem 3.4.2]) sending points of $\widetilde{\operatorname{\mathscr{A}}}_{0}$ to corresponding conjugate filtration. Let
@@
[…]
By previous method of constructing local model diagram of splitting models, there is a local model diagram:
[…]
@@
Note that for each $1\leq i\leq r$, by construction we have in fact a universal $\operatorname{\mathcal{G}}_{0,i}^{\mathrm{spl}}$-zip of type $\mu_{i}$ over $\mathscr{A}^{\mathrm{spl}}_{0}$, thus a morphism
@@
The universal $\mathcal{G}_{0}^{\mathrm{spl}}$-zip of type $\mu$ over $\mathscr{A}_{0}^{\mathrm{spl}}$ is essentially equivalent to the product of all the $\operatorname{\mathcal{G}}_{0,i}^{\mathrm{spl}}$-zip of type $\mu_{i}$.

### v3

@@ 2.2. Smooth Pappas-Rapoport splitting models
By abuse of notation, we still denote by $\mathcal{V}$ the integral automorphic vector bundle over $\operatorname{\mathscr{A}}^{\mathrm{spl}}$. We discuss the example of the standard representation $\Lambda^{\mathrm{spl}}$. Let $(A^{\textrm{univ}},\lambda,\iota)$ be the universal abelian scheme over $\operatorname{\mathscr{A}}^{\mathrm{spl}}$ and $\mathcal{H}=H^{1}_{\operatorname{dR}}(A^{\textrm{univ}}/\operatorname{\mathscr{A}}^{\mathrm{spl}})$. Then as before we have decompositions $\mathcal{H}=\bigoplus_{i}\mathcal{H}_{i}^{m_{i}},\mathcal{H}_{i}=\bigoplus_{j}\mathcal{H}_{i,j}$. We have also the universal splitting structure $\underline{\operatorname{\mathscr{F}}_{\bullet}}$ on $\mathcal{H}$. For each $i,j,l$, we define
@@ 3.1. $F$-zips and $G$-zips
  • $C^{\bullet}=(C^{i})_{i\in\mathbb{Z}}$ is a descending filtration on $M$ and each $C^{i}$ is locally free $\operatorname{\mathcal{O}}_{S}$-module;
@@
  • $\varphi_{\bullet}=(\varphi_{i})_{i\in\mathbb{Z}}$ and for each $i$, $\varphi_{i}:C^{i}/C^{i+1}\to D_{i}/D_{i-1}$ is a $\sigma$-linear map whose linearization
@@
Next we briefly review the notion of $G$-zips and the theory of $G$-zip stacks ([53, 54]) for later use in this section. Let $G$ be a connected reductive group over $\mathbb{F}_{p}$, and $\chi$ a cocharacter of $G$ defined over a finite field $\kappa|\mathbb{F}_{p}$. Let $P_{+}$ (resp. $P_{-}$) be the parabolic subgroup of $G_{\kappa}$ such that its Lie algebra is the sum of spaces with non-negative weights (resp. non-negative weights) in $\operatorname{Lie}(G_{\kappa})$ under $\operatorname{Ad}\circ\chi$. We will also write $U_{+}$ (resp. $U_{-}$) for the unipotent radical of $P_{+}$ (resp. $P_{-}$). Let $L$ be the common Levi subgroup of $P_{+}$ and $P_{-}$.
@@ A $G$-zip of type $\chi$ over $S$ is a tuple $\underline{I}=(I,I_{+},I_{-},\iota)$ consisting of
${E_{G,\chi}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${P_{-}^{(p)}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${P_{+}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${L\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\operatorname{Frob}_{p}}$${L^{(p)}}$
@@
We have an isomorphism of algebraic stacks $E_{G,\chi}\backslash G_{\kappa}]\simeq{G}\text{-}\mathrm{Zip}^{\chi}_{\kappa}$.
@@
    Denote by $|[E_{G,\chi}\backslash G_{\kappa}]\otimes k|$ the topological space of $E_{G,\chi}\backslash G_{\kappa}]\otimes k$, and $^{J}W$ the topological space induced by the partial order $\preceq$. Then the association $w\mapsto G_{w}$ induces a homeomorphism ${J}W\simeq|[E_{G,\chi}\backslash G_{\kappa}]|$.
@@
    $\underline{\mathscr{F}_{\bullet}}=(\mathscr{F}_{i,j}^{l})$ is a splitting structure of $\operatorname{\mathcal{O}}_{F}\otimes_{\mathbb{Z}_{p}}\operatorname{\mathcal{O}}_{S}$-module $\omega_{A/S}$. This means: if we write $\mathcal{H}=\mathcal{H}_{\operatorname{dR}}^{1}(A/S)$, it is an $\operatorname{\mathcal{O}}_{F}\otimes\operatorname{\mathcal{O}}_{S}$-module, hence has a decomposition
@@
Note that each $\mathcal{M}_{i,j}^{l}$ is a locally free $\operatorname{\mathcal{O}}_{S}$-module of rank $2d_{i}$. As in the description below Definition 2.1, in fact we have some reduced description for $\mathcal{M}_{i,j}^{l}$ according to the type of $i$, which we leave to the reader. Now we write
@@
    which is an $\operatorname{\mathcal{O}}_{S}$-linear morphism. On the other side, the injection $\varepsilon_{i}^{-1}\mathscr{F}_{i,j}^{l-2}\hookrightarrow\varepsilon_{i}^{-1}\mathscr{F}_{i,j}^{l-1}$ induces an $\operatorname{\mathcal{O}}_{S}$-linear morphism
@@
${{\mathcal{M}_{i}^{1}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{F_{i}^{2}}$${{\mathcal{M}_{i}^{2}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{V_{i}^{2}}$$\scriptstyle{F_{i}^{3}}\textstyle{\cdots\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{V_{i}^{3}}$$\scriptstyle{F_{i}^{e_{i}}}$${{\mathcal{M}_{i}^{e_{i}}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces.}$$\scriptstyle{V_{i}^{e_{i}}}$
@@
${{\mathcal{M}_{i}^{e_{i}}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{F_{i}^{1}}$${{\mathcal{M}_{i}^{1}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces.}$$\scriptstyle{V_{i}^{1}}$
@@
There is a canonical isomorphism $g_{i}:\operatorname{\mathcal{M}}^{1}_{i}\to\operatorname{\mathcal{M}}^{1}_{i}$ sending $\operatorname{Ker}(V_{i}^{1})$ to $\operatorname{Im}(F_{i}^{1})$.
@@
For each $j$, fix an isomorphism $\operatorname{\mathcal{H}}_{i,j}\simeq R[\varepsilon_{i}]^{2d_{i}}$ and lift it to $\widetilde{\operatorname{\mathcal{H}}}_{i,j}=R[[t]]^{2d_{i}}$ such that $\widetilde{\operatorname{\mathcal{H}}}_{i,j}/(t^{e_{i}})=\operatorname{\mathcal{H}}_{i,j}$. For each $\operatorname{\mathscr{F}}_{i,j}^{l}\subset\operatorname{\mathcal{H}}_{i,j}$, let $\widetilde{\operatorname{\mathscr{F}}}_{i,j}^{l}$ be its lifting in $\widetilde{\operatorname{\mathcal{H}}}_{i,j}$. One can lift the Frobenius and Verschiebung morphisms of $\operatorname{\mathcal{H}}_{i}$ to injections of $\widetilde{\operatorname{\mathcal{H}}}_{i}$, such that
@@
${t^{e_{i}-1}\widetilde{\operatorname{\mathcal{H}}}_{i,j}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\operatorname{Ver}}\scriptstyle{\simeq}$${t^{e_{i}-1}\widetilde{\operatorname{\mathscr{F}}}_{i,j-1}^{e_{i}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\operatorname{Frob}}\scriptstyle{\simeq}$${t^{e_{i}-1}t^{e_{i}}\widetilde{\operatorname{\mathcal{H}}}_{i,j}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\cdot t^{-e_{i}}}\scriptstyle{\simeq}$${t^{e_{i}-1}\widetilde{\operatorname{\mathcal{H}}}_{i,j}.}$
Such an isomorphism sends the lifting of $\operatorname{Ker}(V_{i,j}^{1})$ to the lifting of $\operatorname{Im}(F_{i,j}^{1})$. After modding out by $(t^{e_{i}})$ of the above isomorphism, we get an isomorphism
@@
Such a datum gives the parabolic subgroups $P_{+},P_{-}$ of $\operatorname{\mathcal{G}}_{\kappa}^{\mathrm{spl}}$ and corresponding unipotent groups $U_{+},U_{-}$, and common Levi subgroup $L=P_{+}\cap P_{-}$ as in previous subsection. The cocharacter $\mu$ induces a standard $F$-zip
@@
[…]
@@
${\widetilde{\operatorname{\mathscr{A}}}^{\mathrm{spl}}_{0}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi}$$\scriptstyle{q}$${\operatorname{\mathscr{A}}^{\mathrm{spl}}_{0}}$${M^{\mathrm{loc}},}$
where $M^{\mathrm{loc}}$ is the special fiber of $\mathbb{M}^{\mathrm{loc}}(\mathcal{G}^{\mathrm{spl}},\mu)$, $q$ sends $(x=(\underline{A},\underline{\operatorname{\mathscr{F}}}_{\bullet}),\tau:\operatorname{\mathcal{M}}_{x}\simeq\Lambda^{\mathrm{spl}}_{0})$ to $\tau(\operatorname{\mathcal{M}}_{x}\supset C_{x})\in M^{\mathrm{loc}}=\operatorname{\mathcal{G}}_{0}^{\mathrm{spl}}/P_{+}$. By the same arguments as in [59, §3.4], we have
@@
where $M^{\mathrm{loc},c}(\operatorname{\mathcal{G}}_{0},\mu)$ classifies the conjugate filtration of $\operatorname{\mathscr{A}}_{0}$ defined in [59] and $q^{c}$ is a smooth morphism (same as [59, Theorem 3.4.2]) sending points of $\widetilde{\operatorname{\mathscr{A}}}_{0}$ to corresponding conjugate filtration. Let
@@
${\widetilde{\operatorname{\mathscr{A}}}^{\mathrm{spl}}_{0}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi}$$\scriptstyle{q^{c}}$${\operatorname{\mathscr{A}}^{\mathrm{spl}}_{0}}$${M^{\mathrm{spl},c}(\operatorname{\mathcal{G}}_{0},\mu).}$
By the previous method of constructing local model diagram of splitting models, there is a local model diagram:
${\widetilde{\operatorname{\mathscr{A}}}^{\mathrm{spl}}_{0}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi}$$\scriptstyle{q^{c}}$${\operatorname{\mathscr{A}}^{\mathrm{spl}}_{0}}$${M^{\mathrm{loc},c}}$
@@
Note that for each $1\leq i\leq r$, by construction we have in fact a universal $\operatorname{\mathcal{G}}_{i,0}^{\mathrm{spl}}$-zip of type $\mu_{i}$ over $\mathscr{A}^{\mathrm{spl}}_{0}$, thus a morphism
@@
Via the inclusion $\mathcal{G}_{0}^{\mathrm{spl}}\subset\prod_{i}\operatorname{\mathcal{G}}_{i,0}^{\mathrm{spl}}$, the universal $\mathcal{G}_{0}^{\mathrm{spl}}$-zip of type $\mu$ over $\mathscr{A}_{0}^{\mathrm{spl}}$ induces a $\prod_{i}\operatorname{\mathcal{G}}_{i,0}^{\mathrm{spl}}$-zip of type $\prod_{i}\mu_{i}$, which is isomorphic to the product of all the $\operatorname{\mathcal{G}}_{i,0}^{\mathrm{spl}}$-zip of type $\mu_{i}$.